# Distributional Lusternik–Schnirelmann category of manifolds

### v2

@@ Abstract.
Recently, a new homotopy invariant of metric spaces, called the distributional LS-category, was defined, which provides a lower bound to the classical LS-category. In this paper, we obtain several sufficient conditions for the distributional LS-category ($\operatorname{\mathsf{dcat}}$) of a closed manifold to be maximum, i.e., equal to its classical LS-category ($\operatorname{\mathsf{cat}}$). These give us many new computations of $\operatorname{\mathsf{dcat}}$, especially for some essential manifolds and (generalized) connected sums. In the process, we also determine the $\operatorname{\mathsf{dcat}}$ of closed $3$-manifolds having torsion-free fundamental groups and some closed geometrically decomposable $4$-manifolds. Finally, we extend some of our results to closed Alexandrov spaces with curvature bounded below and discuss their $\operatorname{\mathsf{cat}}$ and $\operatorname{\mathsf{dcat}}$ in dimension $3$.
@@
Aspherical space, distributional LS-category, connected sum, essential manifold, LS-category, Alexandrov space, distributional topological complexity
@@ 1. Introduction
Recently, in a joint work with Dranishnikov [DJ], we defined a new homotopy invariant of spaces, called the distributional LS-category, denoted $\operatorname{\mathsf{dcat}}(X)$ for a space $X$ (see Section 2.A for its motivation). In general, $\operatorname{\mathsf{dcat}}(X)\leq\operatorname{\mathsf{cat}}(X)$. The invariant $\operatorname{\mathsf{dcat}}$ satisfies most of the nice properties as $\operatorname{\mathsf{cat}}$. However, $\operatorname{\mathsf{dcat}}$ is a different notion than $\operatorname{\mathsf{cat}}$ and has various contrasting properties as well (see Section 2.B for the details). Unlike the case of the classical LS-category, the distributional LS-category is known only for a few classes of spaces [DJ][KW][Dr2], and there are a limited number of tools available for its computation.
The main objective of this paper is to compute $\operatorname{\mathsf{dcat}}$ for various new classes of spaces by highlighting more similarities between the notions of $\operatorname{\mathsf{cat}}$ and $\operatorname{\mathsf{dcat}}$. In this direction, we present various sufficient conditions for the distributional LS-category of a closed manifold to be equal to its classical LS-category. As a result, we obtain analogs of some classical results on the LS-category of closed manifolds (especially the two results mentioned in Paragraph $2$ above) for the distributional LS-category. Our techniques also enable us to partially extend some of these sufficient conditions to a more general class of closed Alexandrov spaces with curvature bounded below, [BGP][BBI]. This helps us determine the classical LS-category of closed Alexandrov spaces in some cases.
In particular, we determine $\operatorname{\mathsf{dcat}}$ of some essential manifolds, connected sums of closed manifolds, orientable generalized connected sums of compact manifolds along closed submanifolds, and connected sums of closed Alexandrov spaces. These results give us various new computations of $\operatorname{\mathsf{dcat}}$. The unified statement of our main results is as follows.
@@
The distributional LS-category is maximum (equal to the dimension of the space) for each of the following closed manifolds and closed Alexandrov spaces.
  1. (1)
    $M$, where $M$ is a closed essential $n$-manifold such that $\pi_{1}(M)$ is torsion-free and $M$ satisfies the cap property (i)(i)(i) The “cap property” for essential manifolds is defined in Section 3.A. (e.g., if $B\pi_{1}(M)$ is a closed $n$-manifold).
  2. (2)
    $M\#N$, where $M$ is a closed $n$-manifold and $N$ is a closed essential $n$-manifold such that $\pi_{1}(N)$ is torsion-free and $N$ satisfies the cap property (e.g., if $B\pi_{1}(N)$ is a closed $n$-manifold or $N$ is aspherical).
  3. (3)
    $M\#_{L}\hskip 0.56905ptN$, where $M$ and $N$ are closed orientable $n$-manifolds glued (ii)(ii)(ii) The generalized connected sum of closed manifolds is defined in Section 6.A. along $L$ and $N/L$ is a closed orientable aspherical $n$-pseudomanifold.
  4. (4)
    $\mathcal{D}M$, where $M$ is the product of finitely many punctured (iii)(iii)(iii) For us, a puncture of an $n$-manifold is the removal of an open $n$-disc from it, see Section 6.B. closed orientable aspherical manifolds and $\mathcal{D}M$ is the double of $M$.
  5. (5)
    $M\#N$, where $M$ is a closed Alexandrov $n$-space, $N$ is a closed aspherical Alexandrov $n$-manifold, and the connected sum is over regular points (iv)(iv)(iv) Regular points and the connected sum of Alexandrov spaces are defined in Section 7.A..
  6. (6)
    $M\#N$, where $M$ and $N$ are closed orientable Alexandrov $n$-spaces, $N$ is aspherical, and the connected sum is over regular points.
We note that the above statements are proven in dimensions $\geq 3$. A closed $1$-manifold is a circle and a closed $2$-manifold is a surface of genus $\geq 0$. Also, an Alexandrov $n$-space is a manifold for $n\leq 2$ (see Section 7.A). The distributional LS-category of these spaces has been completely determined in [DJ, Section 6.1]. In fact, $\operatorname{\mathsf{dcat}}(M)=\operatorname{\mathsf{cat}}(M)$ if $M$ is a closed $n$-manifold for $n\leq 2$ other than ${\mathbb{R}}P^{2}$ and $\operatorname{\mathsf{dcat}}({\mathbb{R}}P^{2})=1$. So, in this paper, we shall always work in dimensions $n\geq 3$ unless explicitly stated otherwise.
The main ingredient of our proofs is obstruction theory. Basically, we use some primary obstructions (see details in Section 3.A) that turn out to be useful for our purpose due to the works of Knudsen and Weinberger [KW] and Dranishnikov [Dr2].
Significance of Theorem A
Here, we highlight the relevance of each of the individual statements from Theorem A for $\operatorname{\mathsf{dcat}}$ and $\operatorname{\mathsf{cat}}$.
Item ((i)) generalizes the fact from [KW] and [Dr2] that $\operatorname{\mathsf{dcat}}(M)=\dim(M)$ for any closed aspherical $n$-manifold $M$. Item (2) gives a partial connected sum formula for $\operatorname{\mathsf{dcat}}$ analogous to the formula of [DS, Theorem 1] for $\operatorname{\mathsf{cat}}$. Items ((ii)) and ((iii)) determine $\operatorname{\mathsf{dcat}}$ of some orientable generalized connected sums that are not covered in item (2). Items ((iv)) and (6) extend item (2) to some non-manifold cases and provide a partial connected sum formula for $\operatorname{\mathsf{cat}}$ and $\operatorname{\mathsf{dcat}}$ of closed (possibly non-manifold) Alexandrov spaces. This is also relevant because the connected sum formula for $\operatorname{\mathsf{cat}}$ from [DS] is known only for closed manifolds.
Applications of Theorem A
Here, we highlight the new computations of $\operatorname{\mathsf{dcat}}$ that are obtained as a consequence of the statements of Theorem A.
Item ((i)) computes $\operatorname{\mathsf{dcat}}$ of several symplectically aspherical manifolds that are not aspherical. Item (2) computes $\operatorname{\mathsf{dcat}}$ of connected sums involving the examples considered in item ((i)) and aspherical manifolds whose rational cup lengths are not maximum. It also helps determine $\operatorname{\mathsf{dcat}}$ of closed $3$-manifolds, see Theorem B. Item ((iii)) computes $\operatorname{\mathsf{dcat}}$ of some closed geometrically decomposable $4$-manifolds [Hi] that are not aspherical. Item ((iv)) calculates $\operatorname{\mathsf{dcat}}$ of connected sums of non-manifolds with some examples considered in item (2). Items ((iv)) and (6) also help in estimating $\operatorname{\mathsf{dcat}}$ and $\operatorname{\mathsf{cat}}$ of closed Alexandrov $3$-spaces in some special cases.
In this paper, we also completely determine the distributional LS-category of closed $3$-manifolds whose fundamental groups do not have any non-trivial torsion. This provides an analog of the main result of [GLGA] and [OR] in all but one case.
@@
$\operatorname{\mathsf{dcat}}(M)=\begin{cases}1&\textup{if }\pi_{1}(M)=0\\ 2&\textup{if }\pi_{1}(M)\neq 0\text{ is free}\\ 3&\textup{if }\pi_{1}(M)\text{ is not free but torsion-free}\end{cases}$
Contrary to the case of the classical LS-category, $\operatorname{\mathsf{dcat}}$ of a closed $3$-manifold is not described entirely by its fundamental group (see Section 5.A). Indeed, the problem arises when the fundamental group has torsion.
We use Theorems A (2) and B to prove an analog of Rudyak’s conjecture [R1][R2] for $\operatorname{\mathsf{dcat}}$ of low-dimensional manifolds in various cases.
@@
In Section 2, we revisit the theory of distributional LS-category and recall some facts and notions from algebraic topology. In Section 3, we prove Theorem A ((i)) and see its various applications. In Section 4, we prove Theorem A (2) and use it to obtain many new computations of $\operatorname{\mathsf{dcat}}$. In Section 5, we study $\operatorname{\mathsf{dcat}}$ of low-dimensional manifolds, prove Theorem B, and apply it to an analog of Rudyak’s conjecture. Section 6 contains the proof of Theorem A ((ii)) and ((iii)), and the application of Theorem A ((iii)) to geometrically decomposable $4$-manifolds. In Section 7, we study Alexandrov spaces, prove Theorem A ((iv)) and (6), and provide some new computations for both $\operatorname{\mathsf{cat}}$ and $\operatorname{\mathsf{dcat}}$. In Section 8, we focus on non-manifold Alexandrov $3$-spaces, discuss some problems in determining their $\operatorname{\mathsf{cat}}$ and $\operatorname{\mathsf{dcat}}$, and describe spaces $X$ for which $\operatorname{\mathsf{cat}}(X)>1$ using their curvature and cohomogeneity. We end this paper in Section 9 which is inspired from [Pa] and can be read independently from the rest of this paper. In this section, we provide some necessary conditions for the distributional topological complexity [DJ] of some closed orientable manifolds to be small, i.e., at most $3$.
@@ 2. Preliminaries
2.A. Motivation for $\operatorname{\mathsf{dcat}}$
Given a pointed topological space $(Z,z_{0})$, let us define $P(Z)=\left\{\phi:[0,1]\to Z\mid\phi(1)=z_{0}\right\}$, and let $p_{Z}:P(Z)\to Z$ be the evaluation fibration defined as $p(\phi)=\phi(0)$ for all $\phi\in P(Z)$. For each $n\geq 1$, we define the $n^{th}$ Ganea space, denoted $G_{n}(Z)$, to be the fiberwise iterated join of $n$ copies of $P(Z)$ along $p_{Z}$, i.e.,
@@ 2.A. Motivation for $\operatorname{\mathsf{dcat}}$
Each element of $G_{n}(Z)$ is a formal ordered linear combination where the terms with weights $\lambda_{i}=0$ are dropped. On this space, we define the $n^{th}$ Ganea fibration $p_{n}^{Z}:G_{n}(Z)\to Z$ such that $p_{n}^{Z}(\sum_{i=1}^{n}\lambda_{i}\phi_{i})=\phi_{i}(0)$ for any $i\leq n$ such that $\lambda_{i}>0$.
@@
2.B. Distributional LS-category
For a metric space $Z$, let $\mathcal{B}(Z)$ denote the set of probability measures on $Z$ and for any $n\geq 1$, let

### v3

# Distributional category of manifolds

@@ Abstract.
Recently, a new homotopy invariant of metric spaces, called the distributional category, was defined, which provides a lower bound to the Lusternik–Schnirelmann (LS) category. In this paper, we obtain several sufficient conditions for the distributional category ($\operatorname{\mathsf{dcat}}$) of a closed manifold to be maximum, i.e., equal to its classical LS-category ($\operatorname{\mathsf{cat}}$). These give us many new computations of $\operatorname{\mathsf{dcat}}$, especially for some essential manifolds and (generalized) connected sums. In the process, we also determine the $\operatorname{\mathsf{dcat}}$ of closed $3$-manifolds having torsion-free fundamental groups and various closed geometrically decomposable $4$-manifolds. Finally, we extend some of our results to closed Alexandrov spaces with curvature bounded below and discuss their $\operatorname{\mathsf{cat}}$ and $\operatorname{\mathsf{dcat}}$ in dimension $3$.
@@
Aspherical space, distributional category, connected sum, essential manifold, Lusternik–Schnirelmann category, Alexandrov space.
@@ 1. Introduction
Recently, in a joint work with Dranishnikov [DJ], we defined a new homotopy invariant of spaces, called the distributional category, denoted $\operatorname{\mathsf{dcat}}(X)$ for a space $X$ (see Section 2.1 for its motivation). In general, $\operatorname{\mathsf{dcat}}(X)\leq\operatorname{\mathsf{cat}}(X)$. The invariant $\operatorname{\mathsf{dcat}}$ satisfies most of the nice properties as $\operatorname{\mathsf{cat}}$. However, $\operatorname{\mathsf{dcat}}$ is a different notion than $\operatorname{\mathsf{cat}}$ and has various contrasting properties as well (see Section 2.2 for the details). Unlike the case of the classical LS-category, the distributional category is known only for a few classes of spaces [DJ][KW][Dr2], and there are a limited number of tools available for its computation.
Since the inception of this nascent invariant whose research is an emerging topic, the following tantalizing question has been around.
1.1 Question.
For what spaces $X$ do we have the equality $\operatorname{\mathsf{dcat}}(X)=\operatorname{\mathsf{cat}}(X)$?
To answer the above question, it is worthwhile to check if the techniques that are typically useful in determining LS-category remain somewhat useful in determining distributional category as well. One such technique is lower bounding the categories of a finite CW complex with its rational cup-length, [DJ]. Another example is the Eilenberg–Ganea technique [EG] of expressing the categories of the classifying space of a torsion-free discrete group in terms of the cohomological dimension of the group, [KW].
As a result of these techniques, some spaces on which the distributional category agrees with the LS-category include closed surfaces (other than ${\mathbb{R}}P^{2}$), complex projective spaces, finite products of spheres, and classifying spaces of torsion-free discrete groups. Spaces on which the two notions of category don’t agree include real projective spaces and classifying spaces of finite groups, [DJ][KW]. These examples make Question 1.1 very interesting.
The main objective of this paper is to answer Question 1.1 by finding general classes of compact metric spaces (in particular, closed manifolds) on which the two notions of category coincide. We do this by presenting various sufficient conditions for the distributional category of a closed manifold to be equal to its classical LS-category. We show that careful use of some obstruction-theoretic methods turns out to be very helpful in computing the distributional category in various cases.
As a result of our sufficient conditions, we obtain analogs of some classical results on the LS-category of closed manifolds (especially the two results mentioned in Paragraph $2$ above) for the distributional category. Our techniques also enable us to partially extend some of these sufficient conditions to a more general class of closed Alexandrov spaces with curvature bounded below, [BGP][BBI]. This helps us determine the classical LS-category of closed Alexandrov spaces in some cases.
Let us describe our main results and methods more precisely.
For closed essential manifolds with torsion-free fundamental groups, we define the notion of cap property in Section 3.1. Using this notion and a characterization of distributional category from Section 2.2, we obtain the following.
@@
Let $M$ be a closed essential $n$-manifold such that $\pi_{1}(M)$ is torsion-free. If $M$ satisfies the cap property, then $\operatorname{\mathsf{dcat}}(M)=\operatorname{\mathsf{cat}}(M)=n$.
To prove Theorem A and other major results, we use some primary obstructions (see details in Section 3.1) that turn out to be quite useful for our purpose due to the works of Knudsen and Weinberger [KW] and Dranishnikov [Dr2].
Using the relationship between the orientation sheaves of a connected sum and its summands, we generalize Theorem A to the following.
@@
Let $M$ be a closed $n$-manifold and $N$ be a closed essential $n$-manifold such that $\pi_{1}(N)$ is torsion-free. If $N$ satisfies the cap property, then we have the equality $\operatorname{\mathsf{dcat}}(M\#N)=\operatorname{\mathsf{cat}}(M\#N)=n$.
As an application of Theorem B, we completely determine the distributional category of closed $3$-manifolds having torsion-free fundamental groups, thereby obtaining an analog of the main result of [GLGA] and [OR] in all but one case.
Theorem C.
@@
$\operatorname{\mathsf{dcat}}(M)=\begin{cases}1&\text{if }\pi_{1}(M)=0\\ 2&\text{if }\pi_{1}(M)\neq 0\text{ is free}\\ 3&\text{if }\pi_{1}(M)\text{ is not free but torsion-free}\end{cases}$
Next, we turn towards some generalized connected sums (or fiber sums) of orientable manifolds. In Section 6.1, we define the notion of the connected sum $M\#_{L}N$ of two closed smooth orientable $n$-manifolds $M$ and $N$ joined along a space $L$ which is obtained using the embeddings of a common closed submanifold $S$ in $M$ and $N$. For these generalized connected sums, we get the following.
Theorem D.
Let $M$ and $N$ be closed smooth orientable $n$-manifolds glued along $L$ (in the sense of Section 6.1) to form $M\#_{L}\hskip 0.56905ptN$. Let $N/L$ be a closed orientable $n$-psuedomanifold. If $N/L$ is aspherical, then $\operatorname{\mathsf{dcat}}(M\#_{L}\hskip 0.56905ptN)=\operatorname{\mathsf{cat}}(M\#_{L}\hskip 0.56905ptN)=n$.
We also look at some other generalized connected sums, namely the doubles of punctured compact manifolds with boundaries. For any $k\geq 1$, let us call the process of the removal of $k$ number of small, disjoint open $n$-discs (or the interiors of closed $n$-discs) from a closed orientable $n$-manifold $M$ a $k$-puncturing of $M$. For the doubles of the finite products of these manifolds, we obtain the following.
Theorem E.
For $1\leq i\leq n$, let $k_{i}$, $r_{i}\geq 1$ and let $M_{i}$ be a closed orientable aspherical $r_{i}$-manifold. Let $X_{i}$ be obtained from $M_{i}$ after a $k_{i}$-puncturing of $M_{i}$. If $M=\prod_{i=1}^{n}X_{i}$, then $\operatorname{\mathsf{dcat}}(\mathcal{D}M)=\sum_{i=1}^{n}r_{i}\hskip 1.42262pt$ for the double $\mathcal{D}M=M\#_{\partial M}\overline{M}$ of $M$.
Finally, we turn towards the class of closed Alexandrov spaces with curvature bounded below. In Section 7.1, we define the notion of the connected sum of closed Alexandrov $n$-spaces along their regular points as a direct extension of the corresponding notion from [FGGGH] for the case $n=3$. Using techniques similar to those used in the proof of Theorem B, we obtain the following.
Theorem F.
Let $M$ and $N$ be closed Alexandrov $n$-spaces. If $N$ is an aspherical manifold, then $\operatorname{\mathsf{dcat}}(M\#N)=\operatorname{\mathsf{cat}}(M\#N)=n$ for the connected sum $M\#N$ along regular points of $M$ and $N$.
By relaxing the manifold condition on $N$ in Theorem F and instead imposing orientability on the spaces involved, we get the following.
Theorem G.
Let $M$ and $N$ be closed orientable Alexandrov $n$-spaces. If $N$ is aspherical, then $\operatorname{\mathsf{dcat}}(M\#N)=\operatorname{\mathsf{cat}}(M\#N)=n$ for the connected sum $M\#N$ along regular points of $M$ and $N$.
To the best of our knowledge, the classical LS-category of closed Alexandrov $n$-spaces has not been studied in the literature. In this paper, we initiate a study in that direction by proving Theorems F and G using some novel ideas.
We note that each of the above statements is proven in dimensions $\geq 3$. A closed $1$-manifold is a circle and a closed $2$-manifold is a surface of genus $\geq 0$. Also, an Alexandrov $n$-space is a manifold for $n\leq 2$ (see Section 7.1). The distributional category of these spaces has been completely determined in [DJ, Section 6.1]. In fact, $\operatorname{\mathsf{dcat}}(M)=\operatorname{\mathsf{cat}}(M)$ if $M$ is a closed $n$-manifold for $n\leq 2$ other than ${\mathbb{R}}P^{2}$ and $\operatorname{\mathsf{dcat}}({\mathbb{R}}P^{2})=1$. So, in this paper, we shall always work in dimensions $n\geq 3$ unless explicitly stated otherwise.
Significance of our results
Here, we highlight the relevance of each of the above theorems for $\operatorname{\mathsf{dcat}}$ and $\operatorname{\mathsf{cat}}$.
Theorem A generalizes the fact from [KW] and [Dr2] that $\operatorname{\mathsf{dcat}}(M)=\dim(M)$ for any closed aspherical $n$-manifold $M$. Theorem B gives a partial connected sum formula for $\operatorname{\mathsf{dcat}}$ analogous to the formula of [DS, Theorem 1] for $\operatorname{\mathsf{cat}}$. Theorems D and E determine $\operatorname{\mathsf{dcat}}$ of some orientable generalized connected sums that are not covered in Theorem B. Theorems F and G extend Theorem B to some non-manifold cases and provide a partial connected sum formula for $\operatorname{\mathsf{cat}}$ and $\operatorname{\mathsf{dcat}}$ of (possibly non-manifold) closed Alexandrov spaces. This is also relevant because the connected sum formula for $\operatorname{\mathsf{cat}}$ from [DS] is not known for closed Alexandrov spaces.
Applications of our results
Here, we highlight the new computations of $\operatorname{\mathsf{dcat}}$ that are obtained as a consequence of the above theorems.
Theorem A computes $\operatorname{\mathsf{dcat}}$ of several symplectically aspherical manifolds that are not aspherical. Theorem B computes $\operatorname{\mathsf{dcat}}$ of connected sums involving the examples considered in Theorem A and aspherical manifolds whose rational cup-lengths are not maximum. It also helps determine $\operatorname{\mathsf{dcat}}$ of closed $3$-manifolds, see Theorem C. Theorem C proves an analog of Rudyak’s conjecture [R1][R2] for $\operatorname{\mathsf{dcat}}$ of low-dimensional manifolds in various cases. Theorem E computes $\operatorname{\mathsf{dcat}}$ and $\operatorname{\mathsf{cat}}$ of some closed geometrically decomposable $4$-manifolds [Hi] that are not aspherical. The computation of the classical LS-category of these spaces seems missing from the literature, so Theorem E indeed produces some new computations. Theorem F calculates $\operatorname{\mathsf{dcat}}$ of connected sums of non-manifolds with some examples considered in Theorem B. Theorems F and G also help in estimating $\operatorname{\mathsf{dcat}}$ and $\operatorname{\mathsf{cat}}$ of closed Alexandrov $3$-spaces in some special cases.
Lastly, we note that distributional category ($\operatorname{\mathsf{dcat}}$) is a fairly new numerical invariant; indeed, many aspects of its behavior remain to be understood fully. We hope that the results of this paper, which suggest a close relationship of $\operatorname{\mathsf{dcat}}$ with other classical invariants and compute $\operatorname{\mathsf{dcat}}$ of various manifolds using some novel techniques, will stimulate interest in this invariant and further research activity.
@@
In Section 2, we revisit the theory of distributional category and recall some facts and notions from algebraic topology. In Section 3, we prove Theorem A and see its various applications. In Section 4, we prove Theorem B and use it to obtain many new computations of $\operatorname{\mathsf{dcat}}$. In Section 5, we study $\operatorname{\mathsf{dcat}}$ of low-dimensional manifolds, prove Theorem C, and apply it to an analog of Rudyak’s conjecture for distributional category. Section 6 contains the proof of Theorems D and E, and the application of Theorem E to closed geometrically decomposable $4$-manifolds. In Section 7, we study closed Alexandrov spaces, prove Theorems F and G, and provide some new computations for both $\operatorname{\mathsf{cat}}$ and $\operatorname{\mathsf{dcat}}$. We end this paper in Section 8, where we focus on non-manifold Alexandrov $3$-spaces, discuss some problems in determining their $\operatorname{\mathsf{cat}}$ and $\operatorname{\mathsf{dcat}}$, and describe spaces $X$ for which $\operatorname{\mathsf{cat}}(X)>1$ using their curvature and cohomogeneity.
Notations and conventions
@@ 2. Preliminaries
2.1. Motivation for $\operatorname{\mathsf{dcat}}$
Given a pointed topological space $(Z,z_{0})$, let us define $P(Z)=\left\{\phi:[0,1]\to Z\mid\phi(1)=z_{0}\right\}$, and let $p_{Z}:P(Z)\to Z$ be the evaluation fibration defined as $p(\phi)=\phi(0)$ for all $\phi\in P(Z)$. For each $n\geq 1$, we define the $n$-th Ganea space, denoted $G_{n}(Z)$, to be the fiberwise iterated join of $n$ copies of $P(Z)$ along $p_{Z}$, i.e.,
@@ 2.1. Motivation for $\operatorname{\mathsf{dcat}}$
Each element of $G_{n}(Z)$ is a formal ordered linear combination where the terms with weights $\lambda_{i}=0$ are dropped. On this space, we define the $n$-th Ganea fibration $p_{n}^{Z}:G_{n}(Z)\to Z$ such that $p_{n}^{Z}(\sum_{i=1}^{n}\lambda_{i}\phi_{i})=\phi_{i}(0)$ for any $i\leq n$ such that $\lambda_{i}>0$.
@@
2.2. Distributional category
For a metric space $Z$, let $\mathcal{B}(Z)$ denote the set of probability measures $\mu=\sum\lambda_{i}\delta_{x_{i}}$ on $Z$, where $\sum\lambda_{i}=1$, $\lambda_{i}\geq 0$, and $\delta_{x_{i}}$ is the point Dirac measure at $x_{i}\in Z$. For any $n\geq 1$, let